OpenStudy (anonymous):

Find the solution for the differential equation (dP/dt)=(0.8-0.2*t)*P When P(0)=1000, Thanks in advance

OpenStudy (dumbcow):

first group or separate the variables dP/P = .8 -.2t dt integrate both sides ln(P) = .8t -.1t^2 +C take e raised to both sides to cancel natural log P(t) = k*e^(.8t-.1t^2), where k is some constant Now apply initial conditions: P(0) = 1000 1000 = k*e^0 --> k = 1000 P(t) = 1000*e^(.8t-.1t^2)

OpenStudy (anonymous):

Fuuark, thanks a lot man :)! Can't believe I didn't get it

OpenStudy (anonymous):

how did u do the elimination of log?

OpenStudy (dumbcow):

e^ln(x) = x

OpenStudy (anonymous):

ok,thanks